A certain company pays dividends based on the overall state of the economy. You think that the economy has three possible states: boom, level, and slump. If the economy is booming, you will receive a $10 dividend. A normal economy will result in a $5 dividend, and a slumping economy will only pay a $2 dividend. Assuming that each state of the economy has a one in three chance of occurring, what is the expected value of your dividend? Select one: A. $5.67 B. $3.30 C. $15.00 D. $17.00
Answer #1

Probability of boom = PBoom = 1/3

Probability of Level = PLevel = 1/3

Probability of Slump = PSlump = 1/3

Dividend in boom = DBoom = $10

Dividend in Level = DLevel = $5

Dividend in Slump = DSlump = $2

Expected Return = PBoom*DBoom + PLevel*DLevel + PSlump*DSlump = 10*(1/3) + 5*(1/3) + 2*(1/3) = $5.67